# Information Loss under Coarse-Grained Partitions: A Discrete Framework for AI Ethics

### v6

@@ 6 Conclusion
This study has established a rigorous mathematical foundation for Coarse Ethics (CE). By introducing the framework of CGPs, we demonstrated that coarse evaluations can be systematically analyzed rather than being treated as vague or arbitrary assessments. A coarse set is defined by constructing a coarse-grained partition over a totally ordered set, and its mathematical properties have been formalized through the definitions, propositions, and theorems in this paper. The key conclusions drawn from this study are as follows:
  1. 1.
    Coarse evaluation methods, as described in CE, can be formulated with mathematical precision, providing a structured approach to assessments.
  2. 2.
    Different coarse evaluation methods can be analyzed within the framework of CGPs. In particular, each evaluation method can be represented as a probability distribution, allowing for probabilistic and statistical comparison.
  3. 3.
    A quantitative approach to evaluating information loss in coarse assessments is provided by Kullback-Leibler Divergence, which measures the deviation of a coarse-grained distribution from the original evaluation.
As AI systems become more complex, the need for interpretable and adaptable explanations grows. Current XAI techniques, such as LIME and SHAP, often provide explanations at a fixed granularity, which may not align with user-specific cognitive requirements. CGPs offer a structured way to dynamically adjust the level of detail in AI explanations, ensuring that different users receive information at an appropriate level of abstraction. This approach could enhance AI-assisted decision-making in critical fields such as medicine, finance, and law, where balancing interpretability and complexity is essential.
This study presents an initial exploration of CGPs, but several limitations remain. Our analysis is restricted to totally ordered sets and assumes a covering condition, meaning that every element of the original set is included within the coarse-grained structure. This differs from previous research, which allows for the relaxation of coverage constraints. In other words, our approach focuses on cases where a coarse evaluation fully encompasses the set of evaluated entities. Additionally, our discussion has been primarily framed within the context of educational assessments, particularly school testing. Future research should investigate the implications of relaxing the covering condition and explore applications in broader ethical and legal decision-making frameworks.

### v7

@@ 6 Conclusion
This paper has provided a rigorous mathematical framework for analyzing coarse-grained evaluation in the spirit of Coarse Ethics (CE). By introducing coarse-grained partitions (CGPs) on finite discrete totally ordered sets, we showed that coarse evaluations need not be treated as vague or arbitrary, but can instead be represented and analyzed in a precise mathematical form.
Our main contribution is threefold. First, we formalized coarse-grained evaluation as a mapping from a fine-grained score scale to a coarser categorical scale. Second, we introduced categorical unification (CU) and thereby defined a KL-based measure, $D_{\mathrm{KL\text{-}CU}}$, of how far the original fine-grained distribution departs from a canonical fine-scale reconstruction induced by the coarse representation under minimal assumptions. Third, we proved that this divergence vanishes only under a highly restrictive condition: within each grain, the original fine-grained distribution must already be uniform. Hence, zero information loss in the sense of $D_{\mathrm{KL\text{-}CU}}$ is not a realistic benchmark for ordinary evaluative practice, but rather a highly exceptional limiting case.
These results also lead naturally to the problem of optimization. Since multiple coarse-grained evaluations may be compatible with the same fine-grained evaluation, the issue is not merely whether coarse-graining is possible, but how it should be designed. Our analysis suggests that minimizing KL-based information loss provides a principled baseline objective, although in real settings this objective may need to be supplemented or replaced by operational, normative, or decision-theoretic considerations. In this sense, the framework developed here should be understood not as a universal rule for categorization, but as a formal basis for comparing alternative design objectives.
Finally, the framework is also relevant to explainable AI. In AI-assisted settings, a system may hold a fine-grained internal evaluation while needing to present a coarser representation that is intelligible to human users. This can be understood as a transformation from a fine-grained evaluation to a coarse-grained one, and our framework provides a natural way to study the trade-off between preserving critical information and reducing coarsening costs such as cognitive burden and explanatory complexity. Future work should extend the present analysis beyond finite discrete ordered settings and investigate richer optimization problems under domain-specific constraints.